OpenStudy (anonymous):

solve 3x+4=5x-3 clear steps @danish071996

OpenStudy (anonymous):

so that will be 2 annd 1

OpenStudy (anonymous):

it will come 5x-3x=4+3

OpenStudy (anonymous):

2x=7 so x=7/2

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

1st step is what zelman said

OpenStudy (anonymous):

that what im confused about , what answer should we get when we do that

OpenStudy (anonymous):

answer is x=7/2

OpenStudy (anonymous):

where do you get the 7 and 2 from ?

OpenStudy (jhannybean):

divide both side of your equation by 2 to isolate the x.

OpenStudy (anonymous):

3x+4=5x-3

OpenStudy (anonymous):

5x-3x=4+3

OpenStudy (jhannybean):

\[2x=7\]\[\frac{2x}{2}=\frac{7}{2}\]

OpenStudy (anonymous):

sorry x=7/2

OpenStudy (anonymous):

@danish071996 why do we swap the original sum around?

OpenStudy (anonymous):

to find x we can take terms having x on 1 side and constants on the other side.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

right after that what do we do

OpenStudy (anonymous):

then it comes 5x-3x=4+3

OpenStudy (anonymous):

yes so that would be 2x=7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

after that?

OpenStudy (anonymous):

divide by 2 on both sides

OpenStudy (anonymous):

7/2 x=3.5

OpenStudy (anonymous):

yep